Please help me solve this equation: log2((x^2)+7)=4 (the first 2 is supposed to be small and at the bottom)
Answer by stanbon(75887)

solve this equation: log2((x^2)+7)=4
Write in exponential form:
x^2+7 = 2^4
x^2+7 = 16
x^2 = 9
x = 3 or x= -3
====================
Cheers,
Stan H.